# Many (most?) column subset selection criteria

### v1

@@ 5.4 No PTAS for Schatten p-norm minimization of the pseudo-inverse
Lemma 5 implies that the pseudo-inverse of a matrix with $k$ unit-norm columns has a minimal Schatten $p$-norm of $k^{1/p}$. We derive an inapproximability threshold of $k^{1/p}\sqrt{(1+\tfrac{1}{4k})}$ for $p>2$.
@@
The setup is the same as in the proof of Theorem 2. Assume that X3C instance is false, and assume that any two columns of $\boldsymbol{C}$ overlap. As in the proof of Theorem 10 conclude that $\|\boldsymbol{C}^{\dagger}\|_{F}\geq\sqrt{k+\tfrac{1}{4}}=k^{1/2}\sqrt{1+\tfrac{1}{4k}}$. For $p>2$, we interpret the Schatten $p$-norms as vector $p$-norms on the singular values, and apply the relation [22, (5.4.21)] between vector $p$-norms,
@@ 6.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{A})\geq k$ and unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$, and parameter $b>0$.
@@ 6.2 No PTAS for two-norm condition number minimization
Lemma 6 implies that the minimal two-norm condition number is equal to 1. We derive an inapproximability threshold of $\sqrt{2}$.
@@
The setup is the same as in the proof of Theorem 2. We derive a lower bound for the condition numbers for the reduced instance of X3C when it is false. Lemma 7 implies that the condition numbers decrease with the removal of columns. Hence the smallest condition numbers are found by identifying two columns for a false instance of X3C.
@@ 6.3 No PTAS for Frobenius norm condition number minimization
Lemma 6 implies that a matrix with $k$ unit-norm columns has a minimal Frobenius norm condition number of $k$. We derive an inapproximability threshold of $\sqrt{k\left(k+\tfrac{1}{4}\right)}$.
@@
From Theorem 10 follows $\|\boldsymbol{C}^{\dagger}\|_{F}\geq\sqrt{k+\tfrac{1}{4}}$. Regardless of overlaps, the columns of $\boldsymbol{C}$ have unit norm, so (10) implies $\|\boldsymbol{C}\|_{F}=\sqrt{k}$. Thus
@@ 6.4 No PTAS for mixed condition number minimization
Lemma 6 implies that a matrix with $k$ unit-norm columns has a mixed condition number with a minimal value of $\sqrt{k}$. We derive an approximability threshold of $\sqrt{\tfrac{3}{2}k}$.
@@
Regardless of overlaps, the columns of $\boldsymbol{C}$ have unit norm, so (10) implies $\|\boldsymbol{C}\|_{F}^{2}=k$. Thus $\kappa_{D}(\boldsymbol{C})^{2}\geq\frac{3}{2}k$.
@@ 7.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{A})\geq k$ and unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$, and parameter $b>0$.
@@
The proof is the same as that of Theorem 3 but uses Lemma 8 instead.
@@ 7.2 No PTAS for stable rank maximization
Lemma 8 implies that, for a matrix with $k$ unit-norm columns, the maximal value of the stable rank equals $k$. We derive an inapproximability threshold of $\tfrac{3}{4}k$.
@@
because, regardless of overlaps, the columns of $\boldsymbol{C}$ have unit norm, so (10) implies $\|\boldsymbol{C}\|_{F}^{2}=k$.
@@ 8.1 Maximal volume and S-optimality
    Thus $\operatorname{\mathrm{vol}}(\boldsymbol{C})\leq 1$, which shows the inequality. Alternatively this also follows from Hadamard’s inequality [22, Corollary 7.8.3] and the fact that the columns have unit two-norm.
@@
    S-optimality maximization: The inequality follows from Hadamard’s inequality [22, Theorem 7.8.1],
@@
    The equality follows from the assumption that $\boldsymbol{C}$ has unit-norm columns, so that $\operatorname{\mathrm{s_{opt}}}(\boldsymbol{S})=\operatorname{\mathrm{vol}}(\boldsymbol{C})^{1/k}$, and from the previous part. A more complicated proof of the equality is presented in [32, Theorem 3.2].
@@ 8.3 Minimal two-norm and Schatten p-norms
    Conversely, if $\|\boldsymbol{C}\|_{2}=1$, then (10) and the singular value ordering imply $\sigma_{1}(\boldsymbol{C})=\dots=\sigma_{k}(\boldsymbol{C})=1$. That is, $\boldsymbol{C}$ has orthonormal columns.
@@
    Schatten $p$-norm for $p>2$: The norm inequalities [22, Equation (5.4.21)] imply
@@
    As for the equality, we use the same idea as in the proof of [16, Proposition 3.1]. If $\boldsymbol{C}$ has orthonormal columns, then all the singular values of $\boldsymbol{C}$ are equal to $1$ and $\|\boldsymbol{C}\|_{(p)}=k^{1/p}$.
    Conversely, if $\|\boldsymbol{C}\|_{(p)}=k^{1/p}$, then (10) implies
@@
Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ and $\|\boldsymbol{C}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq k$. For $1\leq p<2$, [22, (5.4.21)] implies
@@ 8.4 Minimal pseudo-inverse norms
    Two-norm of the pseudo-inverse: From (10) follows
@@
    Conversely, if $\|\boldsymbol{C}^{\dagger}\|_{2}=1$, then the singular value ordering $\sigma_{1}(\boldsymbol{C})\geq\cdots\geq\sigma_{k}(\boldsymbol{C})=1$ and (10) imply $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$. Thus, $\boldsymbol{C}$ has orthonormal columns.
@@
    Conversely, if $\|\boldsymbol{C}^{\dagger}\|_{(p)}=k^{1/p}$, then (10) implies
@@ 8.5 Minimal condition numbers
    Conversely, if $\kappa_{F}(\boldsymbol{C})=k$, then (10) implies $\|\boldsymbol{C}\|_{F}=\sqrt{k}$, hence $\|\boldsymbol{C}^{\dagger}\|_{F}^{2}=\sum_{j=1}^{k}\sigma_{j}(\boldsymbol{C})^{-2}=k$. As in the proof of Lemma 5, one shows $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$. Thus, $\boldsymbol{C}$ has orthonormal columns.
@@
    Mixed condition number: The inequality follows from the definition of $\kappa_{D}(\boldsymbol{C})$. As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then $\kappa_{D}(\boldsymbol{C})=\sqrt{k}$. Conversely, if $\kappa_{D}(\boldsymbol{C})=\sqrt{k}$, then (10) implies $\sigma_{k}(\boldsymbol{C})=1$. The singular value ordering $\sigma_{1}(\boldsymbol{C})\geq\cdots\geq\sigma_{k}(\boldsymbol{C})=1$ and (10) imply $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$. Thus $\boldsymbol{C}$ has orthonormal columns.
@@
    Conversely, if $\kappa_{D,p}(\boldsymbol{C})={k}^{1/p}$, then $\sum_{j=1}^{p}\sigma_{j}(\boldsymbol{C})^{p}=k\sigma_{k}(\boldsymbol{C})^{p}$, so that
    $\sum_{j=1}^{p}(\sigma_{j}(\boldsymbol{C})^{p}-\sigma_{k}(\boldsymbol{C})^{p})=0.$
@@ 8.6 Maximal stable rank
    Hence $\sum_{j=1}^{k}{(\sigma_{1}(\boldsymbol{C})^{2}-\sigma_{j}(\boldsymbol{C})^{2})}=0$. From the singular value ordering (1) follows that this is a sum of non-negative summands, and every summand must be equal to zero. Hence, all singular values are the same. From (10) follows that they must be equal to one, so $\boldsymbol{C}$ has orthonormal columns.
@@
    As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then all $k$ singular values of $\boldsymbol{C}$ are equal to $1$, and $\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})=k$, see also [23, Example 2.1].
@@
    Hence $\sum_{j=1}^{k}{(\sigma_{1}(\boldsymbol{C})^{p}-\sigma_{j}(\boldsymbol{C})^{p})}=0$. From the singular value ordering (1) follows that this is a sum of non-negative summands, and every summand must be equal to zero. Hence, all singular values of $\boldsymbol{C}$ are the same. From (10) follows that they must be equal to one, so $\boldsymbol{C}$ has orthonormal columns.
@@ 9 Expressions for partitioned pseudo-inverses
and $\boldsymbol{P}_{1}$ and $\boldsymbol{P}_{2}$ are orthogonal projectors. Since $\|\boldsymbol{P}_{1}\|_{2}=\|\boldsymbol{P}_{2}\|_{2}=1$, the product inequalities for singular values [22, Problem 7.3.P16] imply
@@
Therefore $\|\boldsymbol{M}_{1}^{\dagger}\|_{F}\geq\|\boldsymbol{C}_{1}^{\dagger}\|_{F}$. This is a special case of [28, Theorem 3.2]. The proof for $\|\boldsymbol{M}_{2}^{\dagger}\|_{F}\geq\|\boldsymbol{C}_{2}^{\dagger}\|_{F}$ is analogous. Thus

### v2

@@ 5.4 No PTAS for Schatten p-norm minimization of the pseudo-inverse
Lemma 5 in Section 8.4 implies that the pseudo-inverse of a matrix with $k$ unit-norm columns has a minimal Schatten $p$-norm of $k^{1/p}$. We derive an inapproximability threshold of $k^{1/p}\sqrt{(1+\tfrac{1}{4k})}$ for $p>2$.
@@
The setup is the same as in the proof of Theorem 2. Assume that X3C instance is false, and assume that any two columns of $\boldsymbol{C}$ overlap. As in the proof of Theorem 10 conclude that $\|\boldsymbol{C}^{\dagger}\|_{F}\geq\sqrt{k+\tfrac{1}{4}}=k^{1/2}\sqrt{1+\tfrac{1}{4k}}$. For $p>2$, we interpret the Schatten $p$-norms as vector $p$-norms on the singular values, and apply the relation [26, (5.4.21)] between vector $p$-norms,
@@ 6.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$; integer $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$; and parameter $b>0$.
@@ 6.2 No PTAS for two-norm condition number minimization
Lemma 6 in Section 8.5 implies that the minimal two-norm condition number is equal to 1. We derive an inapproximability threshold of $\sqrt{2}$.
@@
The setup is the same as in the proof of Theorem 2. We derive a lower bound for the condition numbers for the reduced instance of X3C when it is false. Lemma 7 in Section 8.5 implies that the condition numbers decrease with the removal of columns. Hence the smallest condition numbers are found by identifying two columns for a false instance of X3C.
@@ 6.3 No PTAS for Frobenius norm condition number minimization
Lemma 6 in Section 8.5 implies that a matrix with $k$ unit-norm columns has a minimal Frobenius norm condition number of $k$. We derive an inapproximability threshold of $\sqrt{k\left(k+\tfrac{1}{4}\right)}$.
@@
From Theorem 10 follows $\|\boldsymbol{C}^{\dagger}\|_{F}\geq\sqrt{k+\tfrac{1}{4}}$. Regardless of overlaps, the columns of $\boldsymbol{C}$ have unit norm, so (10) in Section 8.1 implies $\|\boldsymbol{C}\|_{F}=\sqrt{k}$. Thus
@@ 6.4 No PTAS for mixed condition number minimization
Lemma 6 in Section 8.5 implies that a matrix with $k$ unit-norm columns has a mixed condition number with a minimal value of $\sqrt{k}$. We derive an approximability threshold of $\sqrt{\tfrac{3}{2}k}$.
@@
Regardless of overlaps, the columns of $\boldsymbol{C}$ have unit norm, so (10) in Section 8.1 implies $\|\boldsymbol{C}\|_{F}^{2}=k$. Thus $\kappa_{D}(\boldsymbol{C})^{2}\geq\frac{3}{2}k$.
@@ 7.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$; $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$; and parameter $b>0$.
@@
The proof is the same as that of Theorem 3 but uses Lemma 8 in Section 8.6 instead.
@@ 7.2 No PTAS for stable rank maximization
Lemma 8 in Section 8.6 implies that, for a matrix with $k$ unit-norm columns, the maximal value of the stable rank equals $k$. We derive an inapproximability threshold of $\tfrac{3}{4}k$.
@@
because, regardless of overlaps, the columns of $\boldsymbol{C}$ have unit norm, so (10) in Section 8.1 implies $\|\boldsymbol{C}\|_{F}^{2}=k$.
@@ 8.1 Maximal volume and S-optimality
    Thus $\operatorname{\mathrm{vol}}(\boldsymbol{C})\leq 1$, which shows the inequality. Alternatively this also follows from Hadamard’s inequality [26, Corollary 7.8.3] and the fact that the columns have unit two-norm.
@@
    S-optimality maximization: The inequality follows from Hadamard’s inequality [26, Theorem 7.8.1],
@@
    The equality follows from the assumption that $\boldsymbol{C}$ has unit-norm columns, so that $\operatorname{\mathrm{s_{opt}}}(\boldsymbol{S})=\operatorname{\mathrm{vol}}(\boldsymbol{C})^{1/k}$, and from the previous part. A more complicated proof of the equality is presented in [39, Theorem 3.2].
@@ 8.3 Minimal two-norm and Schatten p-norms
    Conversely, if $\|\boldsymbol{C}\|_{2}=1$, then (10) in Section 8.1 and the singular value ordering imply $\sigma_{1}(\boldsymbol{C})=\dots=\sigma_{k}(\boldsymbol{C})=1$. That is, $\boldsymbol{C}$ has orthonormal columns.
@@
    Schatten $p$-norm for $p>2$: The norm inequalities [26, Equation (5.4.21)] imply
@@
    As for the equality, we use the same idea as in the proof of [19, Proposition 3.1]. If $\boldsymbol{C}$ has orthonormal columns, then all the singular values of $\boldsymbol{C}$ are equal to $1$ and $\|\boldsymbol{C}\|_{(p)}=k^{1/p}$.
    Conversely, if $\|\boldsymbol{C}\|_{(p)}=k^{1/p}$, then (10) in Section 8.1 implies
@@
Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ and $\|\boldsymbol{C}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq k$. For $1\leq p<2$, [26, (5.4.21)] implies
@@ 8.4 Minimal pseudo-inverse norms
    Two-norm of the pseudo-inverse: From (10) in Section 8.1 follows
@@
    Conversely, if $\|\boldsymbol{C}^{\dagger}\|_{2}=1$, then the singular value ordering $\sigma_{1}(\boldsymbol{C})\geq\cdots\geq\sigma_{k}(\boldsymbol{C})=1$ and (10) in Section 8.1 imply $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$. Thus, $\boldsymbol{C}$ has orthonormal columns.
@@
    Conversely, if $\|\boldsymbol{C}^{\dagger}\|_{(p)}=k^{1/p}$, then (10) in Section 8.1 implies
@@ 8.5 Minimal condition numbers
    Conversely, if $\kappa_{F}(\boldsymbol{C})=k$, then (10) in Section 8.1 implies $\|\boldsymbol{C}\|_{F}=\sqrt{k}$, hence $\|\boldsymbol{C}^{\dagger}\|_{F}^{2}=\sum_{j=1}^{k}\sigma_{j}(\boldsymbol{C})^{-2}=k$. As in the proof of Lemma 5, one shows $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$. Thus, $\boldsymbol{C}$ has orthonormal columns.
@@
    Mixed condition number: The inequality follows from the definition of $\kappa_{D}(\boldsymbol{C})$. As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then $\kappa_{D}(\boldsymbol{C})=\sqrt{k}$. Conversely, if $\kappa_{D}(\boldsymbol{C})=\sqrt{k}$, then (10) in Section 8.1 implies $\sigma_{k}(\boldsymbol{C})=1$. The singular value ordering $\sigma_{1}(\boldsymbol{C})\geq\cdots\geq\sigma_{k}(\boldsymbol{C})=1$ and (10) imply $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$. Thus $\boldsymbol{C}$ has orthonormal columns.
@@
    Conversely, if $\kappa_{D,p}(\boldsymbol{C})={k}^{1/p}$, then $\sum_{j=1}^{{\color[rgb]{0,0,0}k}}\sigma_{j}(\boldsymbol{C})^{p}=k\sigma_{k}(\boldsymbol{C})^{p}$, so that
    $\sum_{j=1}^{{\color[rgb]{0,0,0}k}}(\sigma_{j}(\boldsymbol{C})^{p}-\sigma_{k}(\boldsymbol{C})^{p})=0.$
@@ 8.6 Maximal stable rank
    Hence $\sum_{j=1}^{k}{(\sigma_{1}(\boldsymbol{C})^{2}-\sigma_{j}(\boldsymbol{C})^{2})}=0$. From the singular value ordering (1) follows that this is a sum of non-negative summands, and every summand must be equal to zero. Hence, all singular values are the same. From (10) in Section 8.1 follows that they must be equal to one, so $\boldsymbol{C}$ has orthonormal columns.
@@
    As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then all $k$ singular values of $\boldsymbol{C}$ are equal to $1$, and $\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})=k$, see also [27, Example 2.1].
@@
    Hence $\sum_{j=1}^{k}{(\sigma_{1}(\boldsymbol{C})^{p}-\sigma_{j}(\boldsymbol{C})^{p})}=0$. From the singular value ordering (1) follows that this is a sum of non-negative summands, and every summand must be equal to zero. Hence, all singular values of $\boldsymbol{C}$ are the same. From (10) in Section 8.1 follows that they must be equal to one, so $\boldsymbol{C}$ has orthonormal columns.
@@ 9 Expressions for partitioned pseudo-inverses
and $\boldsymbol{P}_{1}$ and $\boldsymbol{P}_{2}$ are orthogonal projectors. Since $\|\boldsymbol{P}_{1}\|_{2}=\|\boldsymbol{P}_{2}\|_{2}=1$, the product inequalities for singular values [26, Problem 7.3.P16] imply
@@
Therefore $\|\boldsymbol{M}_{1}^{\dagger}\|_{F}\geq\|\boldsymbol{C}_{1}^{\dagger}\|_{F}$. This is a special case of [32, Theorem 3.2]. The proof for $\|\boldsymbol{M}_{2}^{\dagger}\|_{F}\geq\|\boldsymbol{C}_{2}^{\dagger}\|_{F}$ is analogous. Thus
@@
Acknowledgments
We are most grateful to three anonymous reviewers for their constructive feedback in regard to exposition and literature references.